OpenStudy (anonymous):

1/x+1/y=1/z xy=z, what is the average of x and y?

hartnn (hartnn):

so u have 1/x+1/y=1/z what to do ?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

flying x method?

OpenStudy (anonymous):

that gives me x+y/xy=1/z

hartnn (hartnn):

does this make sense : (x+y)/xy = 1/z ..........

hartnn (hartnn):

u are fast!

OpenStudy (anonymous):

flying x method for addition

OpenStudy (anonymous):

I got that method out of a 1985 SAT book I think...

OpenStudy (anonymous):

I don't know if they till teach it anymore...

hartnn (hartnn):

now u have xy = z, so put it there

OpenStudy (anonymous):

still

OpenStudy (anonymous):

k...

hartnn (hartnn):

(x+y)/z = 1/z what u get ?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

zx+zy/xy

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

got to put movie for my mom

hartnn (hartnn):

can't u just cancel z from both sides ?

OpenStudy (anonymous):

I'm back

OpenStudy (anonymous):

k

OpenStudy (anonymous):

we can cancel x and y

OpenStudy (anonymous):

to get z+z I think

OpenStudy (anonymous):

so

OpenStudy (anonymous):

zx+zy/xy=2z

hartnn (hartnn):

no! \(\large \frac{x+y}{z}=\frac{1}{z} \implies x+y=1\) got this ?

hartnn (hartnn):

multiplying z on both sides..

OpenStudy (anonymous):

so wait a minute

OpenStudy (anonymous):

x+y/xy

OpenStudy (anonymous):

x+y/xy=1/z

hartnn (hartnn):

put xy=z

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

sometimes I forget to convert...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

x+y=1

OpenStudy (anonymous):

so the average of x and y is 1/2

hartnn (hartnn):

so what's average?

OpenStudy (anonymous):

1/2

hartnn (hartnn):

yup :)